OpenStudy (anonymous):

Guess an antiderivative for the integrand function and Validate your guess by differentiation and then evaluate the given definite integral. (a) integral of xe^(x^2)dx, from 0 to 1 (b) inegral of (x dx/ sqrt (1+x^2)), from 2 to 5

OpenStudy (amistre64):

what are your guesses?

OpenStudy (anonymous):

I don't have any. I don't understand where to start.

OpenStudy (amistre64):

you start with an assumption. use the derivative rules that you were taught to consider a function that could possibly derive into these.

OpenStudy (amistre64):

there is only one rule that i can think of that pertains to an "e" and there are about 2 off hand that i can think of that derive to a sqrt on the bottom

OpenStudy (anonymous):

so is the antiderivative of ex^(x^2) just ex^(x^2)?

OpenStudy (amistre64):

well, we would check the assumptions by actually using the derivative rules: one rule is: e^u derives to u' e^u; so lets assume e^(x^2) and take the derivative to see if we need to adjust it with come constant

OpenStudy (anonymous):

so the derivative of e^(x^2) is...2x(e^(x^2)) right?

OpenStudy (amistre64):

correct, and that compares to our setup in that it has that 2 attached to it ... lets use some arbitrary constant k like this:\[\large k~e^{x^2}\to2k~xe^{x^2}\] such that 2k = 1. what does the value of k need to be?

OpenStudy (anonymous):

k= 1/2

OpenStudy (amistre64):

good, then we know the antiderivative of x e^(x^2) = 1/2 e^(x^2)

OpenStudy (anonymous):

wait..but what happened to "x" in from of "e"?

OpenStudy (amistre64):

we are finding derivatives and anti derivatives .... the derivative of 1/2 e^(x^2) is x e^(x^2) therefore, the antiderivative of x e^(x^2) is 1/2 e^(x^2) we are performing an operation on one function to create another function

OpenStudy (anonymous):

Oh! ok i see now

OpenStudy (amistre64):

do you know how to evaluate an integral over an interval?

OpenStudy (anonymous):

umm no

OpenStudy (amistre64):

it pretty simple spose we have two functions so that: f derives to g the integral of g, over an interval a to b can be formulated by using its anti derivative f: f(b)-f(a)

OpenStudy (anonymous):

ok..so then it would be \[\frac{ 1 }{ 2 }e ^{x ^{2}}]\left(\begin{matrix}1 \\ 0\end{matrix}\right)= \frac{ 1 }{ 2 }e ^{1 ^{2}}-\frac{ 1 }{ 2 }e ^{0 ^{2}}\]?

OpenStudy (amistre64):

very good :)

OpenStudy (anonymous):

so would that be 0- 1/2 ?

OpenStudy (anonymous):

which is -1/2

OpenStudy (amistre64):

no\[\frac12e^1-\frac12e^0\] \[\frac12e-\frac12(1)\] \[\frac12e-\frac12\] \[\frac12(e-1)\]

OpenStudy (anonymous):

ok gotcha

OpenStudy (amistre64):

the other one is similar to this one, the key is in finding a suitable anti derivative what would be your guess of a function that derives to look something like: 1/sqrt(x)

OpenStudy (amistre64):

maybe something function that derives to something looking like:\[\frac{u'}{\sqrt{u}}\]